# A Wiener process perspective on local intrinsic dimension estimation methods

### v1

@@ 4 Wiener process perspective on LID estimation
$\Delta\rho_{t}(x,y)=(\psi*\phi_{t}^{d})(x)\Delta_{y}\phi_{t}^{D-d}(y)+\Delta_{x}(\psi*\phi_{t}^{d})(x)\phi_{t}^{D-d}(y)$ (6)
Since $\phi_{t}^{D-d}(0)=(2\pi t)^{(d-D)/2}$, we obtain
$\Delta_{x}(\psi*\phi_{t}^{d})(x)\phi_{t}^{D-d}(0)=(2\pi t)^{(d-D)/2}\Delta_{x}(\psi*\phi_{t}^{d})(x),$ (7)
which corresponds to the second term of (5).
To derive the first term, we note that a direct computation yields the following.
$\Delta_{y}\phi_{t}^{D-d}(y)=\phi_{t}^{D-d}(y)\left(\frac{\left\lVert{y}\right\rVert^{2}}{t^{2}}+\frac{d-D}{t}\right).$ (8)
Evaluating this at $y=0$ provides the desired result. ∎
@@
5 From Wiener process to LID estimation in LIDL
We are going to use results from the last section to analyze how the first family of algorithms [Tempczyk et al., 2022, Kamkari et al., 2024b] behaves for some particular cases. Because the main contribution from authors of Kamkari et al. [2024b] is an substantial improvement on the side of density estimation, from the theoretical perspective when dealing with perfect density estimators, both algorithms estimate the same quantity and give the same results. This is the reason from now on we will be analyzing LIDL because we want to analyze the aspects of those implementations that do not depend on the problems with density estimation itself.
Given a point $x\in S$ and a set of times $t_{1},\dots,t_{n}$, LIDL estimates the linear regression coefficient $\alpha$ of the set of points $(\log\delta_{i},\log\rho_{t_{i}}(x))$, where $\delta_{i}=\sqrt{t_{i}}$. It is proved in Tempczyk et al. [2022] that
$\log\rho_{t}(x)=(d-D)\log\sqrt{t}+O(1),$ (11)
This procedure can be seen as approximating the asymptotic slope of the parametric curve $(\log\sqrt{t},\log\rho_{t}(x))$, or in other words, the graph of $s\mapsto\log\rho_{e^{2s}}(x)$ for $s\to-\infty$. Another approach would be to study the derivative of this function. Let us define
$\beta_{t}(x)=\frac{2t}{\rho_{t}(x)}\frac{d}{dt}\rho_{t}(x)=\frac{t}{\rho_{t}(x)}\Delta\rho_{t}(x),$ (12)
where the last equality comes from the diffusion equation (1). Moreover, denote the asymptotic slope of the aforementioned curve by
The next proposition will enable us to conclude that indeed these two approaches are equivalent.
Proposition 5.1.
@@ 5 From Wiener process to LID estimation in LIDL
Now, assume condition 3. Since $\lim\log t=-\infty$, for the whole expression to tent to $-\alpha<0$, we need $\lim\log f(t)=\infty$. In this case however we may apply the de l’Hospital rule giving rise to point 4.
@@
$\lim\frac{t(-\alpha t^{-\alpha-1}g(t)+t^{-\alpha}g^{\prime}(t))}{t^{-\alpha}g(t)}=\lim\left(-\alpha+\frac{g^{\prime}(t)}{g(t)}\right)=-\alpha,$ (14)
@@
Now, by putting $f(t)=\rho_{t}(x)$, we can see that the last condition of Proposition 5.1 holds with $\alpha=-\beta(x)/2$. Thus, the equivalent condition 2. yields the following equivalence of both presented approaches.
Corollary 5.2.
@@
Since $\rho_{t}$ is the solution of the diffusion equation on $\mathbb{R}^{D}$ with $C=1/2$, we can replace the time derivative with $\Delta/2$, obtaining by Lemma 4.1
$\beta_{t}(x)=\frac{2t}{\rho_{t}(x)}\frac{d}{dt}\rho_{t}(x)=\frac{t}{\rho_{t}(x)}\Delta\rho_{t}(x)=d-D+\frac{t(2\pi t)^{(d-D)/2}\Delta_{x}(\psi*\phi_{t}^{d})(x)}{\rho_{t}(x)},$ (17)
@@
5.1 Examples
From the theoretical considerations in Tempczyk et al. [2022] it follows, that $\beta(x)=d-D$ if $\psi$ is sufficiently regular and positive near $x$. In other words, in this case
$\lim_{t\to 0^{+}}\frac{\psi*\Delta\phi_{t}^{d}(x)}{\psi*\phi_{t}^{d}(x)}\cdot t=0.$ (18)
In this section, we will try to obtain this conclusion directly in a few special cases by analyzing the behavior of $\beta_{t}(x)$.
The “uniform distribution” on Euclidean space
Clearly, there is no such thing as the uniform distribution on $\mathbb{R}^{d}$. However, from a purely theoretical viewpoint, in our differential equation approach we don’t need the assumption of $\phi$ being a probability density; it could be any function. And since constant functions are usually the simplest examples, we will now investigate what happens if we put $\psi(x)\equiv 1$ on the whole $\mathbb{R}^{d}$ space.
since $\Delta_{x}\psi\equiv 0$. This expression is constant in $t$, and in particular its limit at 00 is $\beta(x)=d-D$. In this case LIDL estimator is not biased for all $t>0$.
Normal distribution
Let us now consider the normal distribution on $\mathbb{R}^{d}$ with covariance matrix $\Sigma=\mathop{\operatorname{diag}}(\sigma_{1}^{2},\dots,\sigma_{d}^{2})$, and denote its density function by $\psi$. The convolution $\psi*\phi_{t}^{d}$ is the density of the normal distribution with covariance matrix $\Sigma+tI$. If we simplify notation be putting $\phi_{i}=\phi_{\sigma_{i}^{2}+t}^{1}$, we get
$\psi*\phi_{t}^{d}(x)=\prod_{i=1}^{d}\phi_{i}(x_{i}).$ (20)
To compute the Laplacian of this convolution, note that
where the first factor does not depend of $x_{i}$, and therefore
leading to
$\beta_{t}(x)=d-D+t\frac{\Delta(\psi*\phi_{t}^{d})(x)}{\psi*\phi_{t}^{d}(x)}=d-D+t\sum_{i=1}^{d}\frac{1}{\phi_{i}(x_{i})}\frac{\partial^{2}\phi_{i}}{\partial x_{i}^{2}}(x_{i})=d-D+t\sum_{i=1}^{d}\frac{x^{2}_{i}-(\sigma_{i}^{2}+t)}{(\sigma_{i}^{2}+t)^{2}}.$ (23)
we have
(a) Example of the bias of a LIDL estimate for different points from $\mathcal{N}(0,1)$ and for different values of $t$. This plot recreates a numerical calculations presented in Figure 4 from Tempczyk et al. [2022].
(b) Plot of a LIDL estimate as a function of time for 3-dimensional Gaussian distribution with $\Sigma=\mathop{\operatorname{diag}}(1,10^{-6},10^{-12})$ and for three different points $x=(0,0,\sigma_{3})$, which represents a different distance from 0.
Figure 1: Deviations from desired LID estimates for Gaussian distributions.
These results express analytically experimental observations from LIDL and FLIPD papers, which we show in Fig. 1. We can also observe, that if we move to the regions of very low probability for a Gaussian, it generates very high positive bias, which may highly overestimate the true LID. Luckily, most of the points in our dataset comes from the region of high probability, but we should be less certain of the estimates for points from low probability regions.
Arbitrary distribution with sufficiently nice density
By this point is should already be clear what a nice density is. We want to be able to use the equality $\Delta_{x}(\psi*\phi_{t}^{d})=\Delta_{x}\psi*\phi_{t}^{d}$. In order to do so, we need $\psi$ to be bounded, twice differentiable, and have bounded first and second-order partial derivatives. Moreover, we will require $\psi$ to have continuous second-order partial derivatives. This is not very restrictive, and many distributions satisfy these properties, including the normal distribution or more generally, mixtures of Gaussians.
In this case, we have
$\lim_{t\to 0^{+}}f*\phi_{t}^{d}(x)=f(x).$ (27)
This gives us, for $x$ such that $\psi(x)>0$,
$\lim_{t\to 0^{+}}\frac{\Delta_{x}\psi*\phi_{t}^{d}}{\psi*\phi_{t}^{d}}\cdot t=\frac{\Delta_{x}\psi(x)}{\psi(x)}\lim_{t\to 0^{+}}t=0,$ (28)
and again $\beta(x)=d-D$. We have proven that in this case $\beta$ yields a correct estimate of dimension, circumventing all the complexities of the proofs in Tempczyk et al. [2022].
It is worth noting, that when $\Delta_{x}\psi=0$, the estimate is accurate. It is the case for the aforementioned ”uniform distribution” on $\mathbb{R}^{d}$, but it is also true if locally the density is a linear function of $x$. In Fig. 1(a) we can observe, that for $x\approx\pm 1$ and small values of $t$, the estimate is accurate.
6 Conclusions
In this work, we have outlined a new perspective for Wiener process based algorithms for LID estimation and shown some results for LIDL, in which we find an analytical description for the phenomena observed in experiments. The future directions from here are: analyzing the uniform distribution on a line segment, adding curvature to this analysis, and an attempt to analyze an NB algorithm in the same way as LIDL.
7 Acknowledgements
Some experiments were performed using the Entropy cluster at the Institute of Informatics, University of Warsaw, funded by NVIDIA, Intel, the Polish National Science Center grant UMO2017/26/E/ST6/00622, and ERC Start- ing Grant TOTAL. The rest of the experiments were performed using the GUŚLARZ 9000 workstation at the Polish Lab for AI (PL4AI).
We want to thank Marek Cygan and Adam Goliński for their overall support during this project.
CRediT Author Statement
Piotr Tempczyk
(35% of the work): Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data Curation, Writing - Original Draft, Writing - Review & Editing, Visualization, Supervision, Project administration.
Adam Kurpisz
(30% of the work): Conceptualization, Methodology, Validation, Formal analysis, Investigation, Writing - Original Draft, Writing - Review & Editing, Visualization, Supervision.
Łukasz Garncarek
(25% of the work): Conceptualization, Methodology, Formal analysis, Investigation, Writing - Original Draft, Writing - Review & Editing.
Dominik Filipiak
(10% of the work): Software, Data Curation.

### v2

@@ Appendix B Laplacian of the diffused density
$\begin{split}\Delta&\rho_{t}(x,y)=\\ &\psi*\phi_{t}^{d}(x)\Delta_{y}\phi_{t}^{D-d}(y)+\Delta_{x}(\psi*\phi_{t}^{d})(x)\phi_{t}^{D-d}(y)\end{split}$ (47)
To derive the first term, we note that a direct computation yields
$\Delta_{y}\phi_{t}^{D-d}(y)=\phi_{t}^{D-d}(y)\left(\frac{\left\lVert{y}\right\rVert^{2}}{t^{2}}+\frac{d-D}{t}\right).\qed$ (48)
As a consequence, by putting $y=0$ and using $\phi_{t}^{D-d}(0)=(2\pi t)^{(d-D)/2}$ we obtain the following.
Corollary B.2.
For $t>0$ and $x\in\mathbb{R}^{d}$ we have
@@
Appendix C Equivalent formulations of LIDL
The following proposition allows us to conclude that the two approaches to LID estimation we discussed are equivalent.
Proposition C.1.
@@ Appendix C Equivalent formulations of LIDL
Now, assume condition 3. Since $\lim\log t=-\infty$, for the whole expression to tend to $-\alpha<0$, we need $\lim\log f(t)=\infty$. In this case however we may apply the de l’Hospital rule giving rise to point 4.
@@
$\begin{split}\lim\frac{t(-\alpha t^{-\alpha-1}g(t)+t^{-\alpha}g^{\prime}(t))}{t^{-\alpha}g(t)}&=\\ &=\lim\left(-\alpha+\frac{g^{\prime}(t)}{g(t)}\right)\\ &=-\alpha,\end{split}$ (52)
@@
Under the notation of Section 5 we have the following two propositions.
Proposition C.2.
@@
Proof.
By using Proposition C.1 with $f(t)=\rho_{t}(x)$, we can see that its last condition holds with $\alpha=-\beta(x)/2$. Thus, the equivalent condition 2. yields the desired equality. ∎
@@
By (8) and Lemma B.1 we obtain
$\begin{split}\beta_{t}(x)&=\frac{t}{\rho_{t}(x)}\Delta\rho_{t}(x)\\ &=d-D+\frac{t(2\pi t)^{(d-D)/2}\Delta_{x}(\psi*\phi_{t}^{d})(x)}{\rho_{t}(x)},\end{split}$ (55)
@@
Appendix D Convex combinations of distributions
Our new approach allows to easily analyze the behavior of LIDL on unions of manifolds. Suppose that $S=\bigcup_{i}S_{i}$ is a finite union of manifolds. The measure $p_{S}$ can be then decomposed into a convex combination of probability measures $p_{i}$ supported on $S_{i}$, namely
$p_{S}=\sum_{i}\lambda_{i}p_{i},$ (56)
where $\lambda_{i}>0$ and $\sum_{i}\lambda_{i}=1$. With each $(S_{i},p_{i})$ we may associate their corresponding diffused density $\rho_{t}^{i}$, reparameterized slope $\beta_{t}^{i}$, and its limit $\beta^{i}$ through equations (3), (8), and (9). Due to bilinearity of convolution, analogous decomposition holds for $\rho_{t}$, namely
$\rho_{t}(x)=\sum_{i}\lambda_{i}\rho_{t}^{i}(x).$ (57)
Using these, we may now attempt to reduce the problem to the study of individual components.
In the following discussion it turns out that the decomposition of $S$ into a union of manifolds, and the fact that $p_{i}$ are supported on the components of this decomposition, are of no importance. The only essential assumption is the convex combination decomposition (56).
Lemma D.1.
A convex combination decomposition $p_{S}=\sum_{i}\lambda_{i}p_{i}$, gives rise to a decomposition of $\beta_{t}(x)$ into a convex combination of $\beta_{t}^{i}(x)$,
$\beta_{t}(x)=\sum_{i}\frac{\lambda_{i}\rho_{t}^{i}(x)}{\rho_{t}(x)}\beta_{t}^{i}(x).$ (58)
In particular, if all the limits below exist, one has
$\beta(x)=\sum_{i}\left(\lim_{t\to 0^{+}}\frac{\lambda_{i}\rho_{t}^{i}(x)}{\rho_{t}(x)}\right)\beta^{i}(x).$ (59)
Proof.
The desired decomposition can be obtained through direct expansion of the definition of $\beta_{t}$. We have
$\begin{split}\beta_{t}(x)&=\sum_{i}\frac{t\lambda_{i}\Delta\rho_{t}^{i}(x)}{\rho_{t}(x)}\\ &=\sum_{i}\frac{\lambda_{i}\rho_{t}^{i}(x)}{\rho_{t}(x)}\cdot\frac{t\Delta\rho_{t}^{i}(x)}{\rho_{t}^{i}(x)}\\ &=\sum_{i}\frac{\lambda_{i}\rho_{t}^{i}(x)}{\rho_{t}(x)}\beta_{t}^{i}(x).\end{split}$ (60)
The second assertion follows by taking the limit. ∎
Now take a closer look at the coefficients $\lambda_{i}\rho_{t}^{i}(x)/\rho_{t}(x)$.
Lemma D.2.
Suppose that $p_{S}=\sum_{i}\lambda_{i}p_{i}$ is a convex combination decomposition, and for some $R>r>0$ there exist $i\neq j$ such that $p_{i}(B(x,R))=0$ and $p_{j}(B(x,r))=C>0$. Then the following estimate holds.
$\frac{\lambda_{i}\rho_{t}^{i}(x)}{\rho_{t}(x)}\leq\frac{\lambda_{i}}{\lambda_{i}+C\lambda_{j}e^{(R^{2}-r^{2})/2t}}.$ (61)
Moreover,
$\lim_{t\to 0^{+}}\frac{\lambda_{i}\rho_{t}^{i}(x)}{\rho_{t}(x)}=0,$ (62)
and $p_{i}$ does not contribute to $\beta(x)$.
Proof.
First, observe that the integral defining $\rho_{t}^{i}(x)$ can be estimated by the supremum of the integrand $\phi_{t}^{D}(x-y)$ over the support of $p_{i}$, contained in $B(x,R)^{c}$, namely
[…]
Therefore, the ratio $\rho_{t}^{j}(x)/\rho_{t}^{i}(x)$ can be bounded from below as
[…]
Combining both these estimates, we end up with
$\begin{split}\frac{\rho_{t}(x)}{\lambda_{i}\rho_{t}^{i}(x)}&=1+\sum_{k\neq i}\frac{\lambda_{k}\rho_{t}^{k}(x)}{\lambda_{i}\rho_{t}^{i}(x)}\\ &\geq 1+\frac{\lambda_{j}\rho_{t}^{j}(x)}{\lambda_{i}\rho_{t}^{i}(x)}\geq 1+\frac{\lambda_{j}}{\lambda_{i}}Ce^{(R^{2}-r^{2})/2t}.\end{split}$ (65)
Inverting and taking the limit yields the desired assertions. ∎